OpenStudy (anonymous):

Solve for x. |2x + 6| - 4 = 20 x = 4 and x = 0 x = -15 and x = 9 x = -4 and x = -2 x = -3 and x = 1 no solutions

OpenStudy (hexagon001):

look at 2 situations... 1) 2x+6 -4 =20 2x +2 =20 2x =20-2 2x =18 x=18/2 x=9 2) -2x-6-4=20 -2x-10=20 -2x =20+10 -2x =20 x= 20/-2 x=-15

OpenStudy (hexagon001):

so x=-15 and 9

OpenStudy (anonymous):

20+10 = 20? ( but yeah ans is correct)

OpenStudy (anonymous):

love you thank you

OpenStudy (hexagon001):

do you get it? for one solution you take what is inside the modulus sign to be all positive and solve it... for the second solution take all in the modulus sign to be negative and then solve it. @luckythebest woopsy that is supposed to read 30

OpenStudy (hexagon001):

@luckythebest and @InsanelyChaotic i typed fast an never realised error.. change the 20 +10 =20 to 20+10=30 and all the subsequent 20s to 30...

OpenStudy (anonymous):

yes, that makes a lot more sense... wow:)